May you please help me with this: y=3x+2
2y-14=-4x

Answer by AnlytcPhil(1810)

y = 3x + 2
2y - 14 = - 4x

Since the first equation is solved for y, and since y is equal
to 3x + 2, we put that in parentheses like this (3x + 2), then
we take the other equation and instead of y put what y equals,
namely (3x + 2).  Then solve for x

         2y - 14 = -4x
  2(3x + 2) - 14 = -4x
     6x + 4 - 14 = -4x 
         6x - 10 = -4x 
         6x + 4x = 10
             10x = 10
               x = 1 
    
Now take the first equation
     
               y = 3x + 2

and substitute (1) for x

               y = 3(1) + 2

then finish the right side

               y = 3 + 2
               y = 5

So the solution is (x,y) = (1,5)

Edwin